OpenStudy (anonymous):

Which expression is equal to (a+b)^2? A. a^2+b^2 B. 2a+2b C. a^2+ab+b^2 D. a^2+2ab+b^2

OpenStudy (mathstudent55):

\((a + b)^2\) \(=(a + b)(a + b) \) Have you learned to multiply two binomials together using FOIL?

OpenStudy (anonymous):

So its D?

OpenStudy (anonymous):

@mathstudent55 ?

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

thanks!!!

OpenStudy (mathstudent55):

\((a + b)(a + b)\) \(=a^2 + ab + ab + b^2\) \(= a^2 + 2ab + b^2\)

OpenStudy (mathstudent55):

You're welcome.